Bean Products Company reported total retained earnings at December 31, Year 2 of $113,500. It provided the following information for Year 1 and Year 2:

Year 2

Year 1

Retained earnings (beginning)

$132,500

$62,500

Revenues

446,000

435,000

Expenses

370,000

312,500

Dividends

?

52,500


How much dividends did the company pay in Year 2?

Select one:

A. $ 95,000

B. $145,000

C. $ 76,000

D. $135,000
Answer #1

Answer: A) $ 95,000

Net income in Year 2 = Revenue -Expenses = 446,00-370,000= 76,000

Devidend Paid in YEar 2 =Retained Earnings (Beginning) + Net income - retianed Earnis closing

= 132,500+76,000-113,500 = 95,000 (Answer)